OpenStudy (anonymous):

5-3+10+(-8) -(-7) solve step by step

OpenStudy (anonymous):

- and + = - + and - = - + and + = + - and - = +

OpenStudy (anonymous):

= 5-3 + 10 -8 + 7 = 2 + 10 -8 + 7 = 12 - 8 + 7 = 4 + 7 = 11

OpenStudy (anonymous):

+(-8) = -8 and -(-7) = +7